# LoRACode: LoRA Adapters for Code Embeddings

### v1

@@ Abstract
Code embeddings are essential for semantic code search; however, current approaches often struggle to capture the precise syntactic and contextual nuances inherent in code. Open-source models such as CodeBERT and UniXcoder exhibit limitations in scalability and efficiency, while high-performing proprietary systems impose substantial computational costs. We introduce a parameter-efficient fine-tuning method based on Low-Rank Adaptation (LoRA) to construct task-specific adapters for code retrieval. Our approach reduces the number of trainable parameters to less than two percent of the base model, enabling rapid fine-tuning on extensive code corpora (2 million samples in 25 minutes on two H100 GPUs). Experiments demonstrate an increase of up to 9.1% in Mean Reciprocal Rank (MRR) for Code2Code search, and up to 86.69% for Text2Code search tasks across multiple programming languages. Distinction in task-wise and language-wise adaptation helps explore the sensitivity of code retrieval for syntactical and linguistic variations.
@@ 1 Introduction
Code embeddings play a crucial role in code search by representing code snippets or programs as vectors in a high-dimensional space. This representation enables computational systems to comprehend the semantic meaning and structural relationships in the code. This technique has become increasingly important due to the growing complexity of software systems and the need for efficient methods to retrieve relevant code from vast codebases.
Code search systems relying on traditional semantic search often struggle to capture the distinct nature of code. Code words have precise meanings, and syntactic variations across languages pose unique challenges for text-based embeddings in code retrieval. Furthermore, since code snippets are context-dependent, they require a broader program context, necessitating specialized models. Although existing models such as CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), UniXcoder (Guo et al., 2022), and StarCoder (Li et al., 2023) are effective, they tend to be smaller in scale and based on BERT-like architectures, limiting their performance in semantic retrieval tasks.
However, existing models, like CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), UniXcoder (Guo et al., 2022), and StarCoder (Li et al., 2023), while effective, are often smaller in scale and based on architectures like BERT, limiting their performance in tasks such as the semantic retrieval of code snippets across text and code-based queries. High-performing models like Voyage-Code 2 (AI, 2024) and OpenAI Embeddings (OpenAI, 2022) offer superior capabilities but are closed-source, expensive, and not designed for code-based retrieval tasks, as they are code-based LLMs rather than specialized code embedding models. Their resource-intensive fine-tuning and lack of scalability for cross-language code search further complicate their use.
@@
2 Background and Related Work
This section provides an overview of related work in the field of code embeddings, covering key models, benchmarks, and techniques. We first discuss various code embedding models, highlighting their strengths and limitations, followed by an exploration of relevant datasets and benchmarks for evaluating code retrieval tasks. The section also delves into techniques such as LoRA and contrastive fine-tuning, which are crucial for enhancing model performance with minimal computational resources.
2.1 Code Embedding Models
CodeBERT (Feng et al., 2020) is a significant advancement that introduces a bimodal pre-trained BERT model for programming and natural languages. While it uses replaced token detection and applies conventional NLP pretraining techniques to source code, it treats code as a simple sequence of tokens, missing important structural information necessary for understanding code semantics. This limitation also affects its scalability and ability to generalize in cross-language tasks due to its reliance on masked language modeling.
GraphCodeBERT (Guo et al., 2021) extends CodeBERT by incorporating a data flow graph to capture structural dependencies in code. This enhancement enables the model to better understand variable usage and control flow, making it more effective for bug detection and code summarization tasks.
UniXcoder (Guo et al., 2022) unifies code representation across multiple modalities, including text, code, and structured representations. By leveraging cross-modal pretraining, UniXcoder achieves state-of-the-art results in tasks like code translation and completion.
@@ 2.1 Code Embedding Models
Utpala et al. (2023) show that previous code embedding models struggle to separate language-specific semantic components from language-agnostic syntactic components during search evaluations. Their paper employs methods to eliminate language-specific information, leading to significant performance improvements in retrieval tasks by focusing on language-agnostic components.
@@ 2.2 Benchmarks and Datasets
XLCost (Zhu et al., 2022) is a benchmark dataset encompassing seven programming languages, focused on aligning code semantics. Its parallel structure allows for effective evaluation of multilingual models, serving as a valuable resource for code reuse and retrieval across different programming environments.
CosQA (Huang et al., 2021) is a question-answering dataset with human-annotated query-code pairs, ensuring strong semantic alignment for text-to-code search tasks. While effective for fine-tuning models, its focus on Python limits its applicability to other programming languages.
@@ 2.4 Embeddings and Contrastive Fine-tuning
Improving text embeddings is a key focus, especially for smaller language models. Techniques like contrastive fine-tuning (Khosla et al., 2021) enhance embeddings by aligning semantically similar text pairs, using contrastive and triplet loss (Chopra et al., 2005; Schroff et al., 2015) from Computer Vision. LoRA further refines these embeddings efficiently by introducing low-rank matrices in attention layers, enabling tailored adjustments for specific datasets or tasks Hu et al. (2021) .
@@
The authors of (Galliamov et al., 2024) mention the usage of PEFT methods like LoRA and Prompt Tuning but do not document adequate information on the implementation or the analysis of their low-rank adaptation approaches.
Liu et al. (2024) highlight advancements in code retrieval using instruction-tuned large language models (LLMs) and evaluates performance with the NDCG@10 metric across eight tasks. It points out the lack of large, open-source embedding models for code retrieval, as most are proprietary. CodeXEmbed outperforms Voyage Code (AI, 2024) on the CoIR benchmark (Li et al., 2024) by 20% and introduces various retrieval models. However, its large model reliance incurs high storage and computation costs. Our work addresses this by utilizing LoRA adapters on smaller code embedding models, significantly reducing resource overhead while maintaining state-of-the-art performance.
@@ 3 Design and Implementation
Each pretrained model – CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), and UniXcoder (Guo et al., 2022) – was fine-tuned using the LoRA configuration. For a detailed explanation of the choice of models, see Appendix A.1. LoRA employs low-rank decomposition to modify only the attention layers while freezing the remaining weights. This approach reduces memory consumption and accelerates training.
@@
In various experiments conducted with StarCoder (Li et al., 2023), we tested different combinations of the Query, Key, and Value layers within the attention mechanism. This was done to analyze the contributions of the early, middle, and outer layers in low-rank decomposition. For more details about the training procedure, please refer to Appendix A.2.
@@ 4.2 Text2Code
To explore the second question highlighted in Section 4, we fine-tune the UniXcoder model for Text2Code retrieval on a single programming language. We use the CosQA (Huang et al., 2021) dataset, as highlighted in Section 4.1. Since CosQA is a question-answering dataset, we experimented with setting the task type for the LoRA Configuration (Houlsby et al., 2019) (Hu et al., 2021) as QUESTION_ANS, which was not feasible for retrieval tasks, since question-answering is a generative task, and would also require the inference using special Roberta (Liu et al., 2019) models encased with question answering, and those suitable for handling parameters like start_position, end_position, etc. The task type was then set to FEATURE_EXTRACTION, the model parameters were frozen, and then the LoRA matrices were trained using Contrastive Trainer. Detailed observations for the same are noted in Appendix A.2.
@@
UniXCoder LoRACode - Combined ($r$=64)
MRR 31.36 36.02
NDCG@10 35.64 40.44
Table 4: Increase in Mean Reciprocal Rank and Normalized Discounted Cumulative Gain @ k=10 for LoRACode combined adapter (rank 64) trained on CosQA dataset (Huang et al., 2021), compared to the UniXCoder base model. LoRACode Combined denotes a single adapter fine-tuned over the 3 base models (UniXcoder, GraphCodeBERT, and CodeBERT), but evaluated on UniXcoder.
@@ 4.3 Language Specific Experimentation
One significant limitation of our approach is the lack of diversity in the data domains used to fine-tune code embedding models. The CodeSearchNet dataset (Husain et al., 2020) mainly extracts code-comment pairs from GitHub, which reflects specific coding practices commonly found in open-source projects. Furthermore, datasets like CodeSearchNet (Husain et al., 2020) and CosQA (Huang et al., 2021) only facilitate text-based retrieval of code snippets. Therefore, the XLCost dataset (Zhu et al., 2022) was essential for conducting our Code2Code search experiments.
@@ C PHP Java C++ C# Javascript Python
Table 6: MRR results for Code2Code search of LoRA models UniXCoder, GraphCodeBERT, and CodeBERT fine-tuned with ranks 32 and 64, compared with base models, evaluated over the XLCost dataset per language. LoRACode Combined denotes a single adapter of given rank fine-tuned over the 3 base models (UniXcoder, GraphCodeBERT and CodeBERT), but evaluated on UniXcoder.
The LoRA adapters consistently outperformed embedding models like GraphCodeBERT and CodeBERT, demonstrating their efficacy in leveraging low-rank adaptations for code retrieval. Different languages report a significant increase in MRR, ranging from 9.1% for C, 7.47% for C++, 6.54% for Java, 5.82% for C#, 4.43% for Javascript, 3.40% for Python, and 3.8% for PHP. Using a LoRA config with rank 32 also substantially increases MRR over the highest-performing code embedding model UniXcoder (Guo et al., 2022). These findings suggest LoRA’s low-rank decomposition plays a crucial role in improving retrieval accuracy, supporting the use of these adapters for multilingual code search. The LoRA config utilizes only 1.83% to 1.85% of the total trainable parameters, leading to memory-efficient and inexpensive fine-tuning. See Appendix A.4 for detailed observations regarding the time taken to generate embeddings for Code2Code search.
@@ Acknowledgments
Appendix A Appendix
A.1 Models Used
We used the code embedding models CodeBERT, GraphCodeBERT, UniXcoder, and StarCoder as base models for parameter efficient fine-tuning. CodeBERT, GraphCodeBERT and UniXcoder were standardly available open-sourced models based on the BERT encoder model, whereas UniXcoder is a pre-trained unified encoder-decoder cross-modal model. RoBERTA based models are also straightforward tokenizers which make for easier fine-tuning. Starcoder is an open-sourced code LLM chosen for its extensive training on over 1 trillion tokens across 80+ programming languages.
CodeBERT demonstrated poor MRR scores because it is based on simple encoder model, whereas the encoder-decoder framework is sub-optimal for autoregressive tasks (Guo et al., 2022). This is why UniXcoder performed well as the base model, and also when adapted with LoRA addends.
A.2 Training Procedure
For fine-tuning a task-specific adapter, we loaded a PEFTConfig of desired rank, set lora_alpha value to be double the rank, set dropout value as 0.1, set the target modules as the query value addends of the Attention layer and occasionally set the task_type for the PEFTConfig to be FEATURE_EXTRACTION. Some details of the implementation procedure and observations are noted as follows:
  • We noticed that the models performed better when set with the FEATURE_EXTRACTION task flag for Text2Code search, but there weren’t much significant improvements for Code2Code search task. This indicates the utility of task type in providing the hidden states that can be used as embeddings for appropriate feature or embedding for the downstream task in question.
  • We tried using a task_type for QUES_ANSWERING for the CosQA dataset (Huang et al., 2021) since it is specifically formatted for question answering tasks, but this would have required a special RobertaTokenizerForQuestionAnswering class during inference rather than the base abstraction RobertaTokenizer needed for MRR evaluation. So the task_type was left at FEATURE_EXTRACTION.
  • Since Starcoder did not have specific query and value modules in its architecture, we discovered the layer-wise modules for each across the 12 layers. We experimented with different combinations of QV targets across the earlier, the middle and the later layers. We found that the aggregate of middle layers query value addends when used as target modules for LoRAConfig, performed better than the other two. The performance was still subpar for StarCoder due to LLMs being inefficient for code retrieval tasks and lack the ability to hold retrieval context during fine-tuning.
A.3 Prevalent Knowledge
We learnt the calculation for last token pooling from similar papers. We thus calculated embeddings by not taking the average of last hidden states, but instead by reducing the masked layers, and only taking an average over the attention heads.
A.4 Diagrams and tables
Table 7 showcases the time taken to generate embeddings in minutes for Code2Code search of the base embedding models when compared with LoRA models. The latency is measured over each programming language’s dataset, for the combined adapter of LoRA rank 64, over UniXCoder as base model.
Base Model PEFT Model
C 0:52 0:44
PHP 4:30 3:44
Java 8:46 9:22
C++ 8:41 9:27
C# 8:37 9:24
Javascript 8:35 8:56
Python 8:24 9:17
Table 7: Time taken to generate embeddings (in MM:SS) for Code2Code search of the base embedding models vs embedding models altered with LoRA adapters. The latency is measured over each programming language’s dataset, for the combined adapter of LoRA rank 64, over UniXCoder as base model.
Figure 2 illustrates the trends in MRR and NDCG for Base models and LoRA-enhanced models across different programming languages. It shows that LoRA consistently improves both retrieval effectiveness (MRR) and ranking relevance (NDCG), with larger performance gains observed in languages with more training data. The visualization highlights the advantage of language-specific fine-tuning with LoRA, demonstrating significant improvements over the baseline models.
Figure 2: Bar Graph showing the trends of MRR and NDCG for Base and LoRA models across different programming languages for Text2Code search.
This Bar Graph shows a substantial increase in both MRR and NDCG, for LoRACode as opposed to the base UniXcoder model. Here LoRACode divided on language-specific adapters perform much better than LoRACode adapters fine-tuned on the aggregation of datasets across all languages, showcasing the importance of tailoring to linguistic diversity.

### v2

@@ Abstract
Code embeddings are essential for semantic code search; however, current approaches often struggle to capture the precise syntactic and contextual nuances inherent in code. Open-source models such as CodeBERT and UniXcoder exhibit limitations in scalability and efficiency, while high-performing proprietary systems impose substantial computational costs. We introduce a parameter-efficient fine-tuning method based on Low-Rank Adaptation (LoRA) to construct task-specific adapters for code retrieval. Our approach reduces the number of trainable parameters to less than two percent of the base model, enabling rapid fine-tuning on extensive code corpora (2 million samples in 25 minutes on two H100 GPUs). Experiments demonstrate an increase of up to 9.1% in Mean Reciprocal Rank (MRR) for Code2Code search, and up to 86.69% for Text2Code search tasks across multiple programming languages. Distinction in task-wise and language-wise adaptation helps explore the sensitivity of code retrieval for syntactical and linguistic variations. To foster research in this area, we make our code and pre-trained models publicly available [^1] Github code repository, Huggingface models trained for LoRACode.
@@ 1 Introduction
Code search based on natural language and code-based queries has long been recognized as a core problem in software engineering, and was in fact, one of the earliest tasks tackled by AI methods. (Sadowski et al., 2015; Allamanis et al., 2018; Xie et al., 2023) denotes the immense applications of probabilistic code search models in recommender systems, inferring code convention, code translation and other similar domains. Over time, the proliferation of open source repositories and question-and-answer forums (e.g., StackOverflow) has only underscored the importance of semantic code search. Modern systems therefore seek rich semantic representations of code; indeed, recent embedding-based retrieval frameworks (for example, Voyage AI’s (AI, 2024) code embedding model and OpenAI’s code embeddings (OpenAI, 2022)) have achieved dramatic gains in search performance, yet stay closed-source and cannot be fine-tuned.
On the other hand, Transformer-based encoders like CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), UniXcoder (Guo et al., 2022), and StarCoder (Li et al., 2023), while effective in learning deep joint representations of code and text, are often smaller in scale and based on architectures like BERT, limiting their performance in tasks such as the semantic retrieval of code snippets across text and code-based queries. These models also fail to generalize over cross-lingual tasks, and hence perform poorly on complex tasks than their variants fine-tuned for specific downstream applications (Xie et al., 2023). With large web-scraped datasets containing about 2 million samples (Husain et al., 2020), fine-tuning also becomes highly resource-intensive.
@@
2 Background & Related Work
This section provides an overview of related work in the field of code search, covering key embedding models, benchmarks, and techniques. We first discuss various code embedding models, highlighting their strengths and limitations, followed by an exploration of relevant datasets and benchmarks for evaluating code retrieval tasks. The section also delves into techniques such as LoRA and contrastive fine-tuning, which are crucial for enhancing model performance with minimal computational resources.
2.1 Code Search and Embedding Models
Earlier research (Sachdev et al., 2018; Allamanis et al., 2015) imagined natural language query-based code search as a statistical modeling problem free from the application of Transformers. Neural Code Search (Sachdev et al., 2018) extracted keywords from function, variable names and comments, and used skip-gram vectorization to create token and document embeddings whose cosine similarity is computed against query embeddings to facilitate search.
While previous work managed to infer merely surface-level features from code tokens and names, CodeBERT (Feng et al., 2020) took a revolutionary dive into learning deep code semantics by introducing a bimodal pre-trained BERT model for programming and natural languages. Trained using Masked Language Modeling and Replaced Token Detection on both source code and natural language, CodeBERT (Feng et al., 2020) leveraged transformer attention mechanisms to model relationships, inviting scope for many more code embedding models possessing the ability to be fine-tuned on a variety of downstream applications. GraphCodeBERT (Guo et al., 2021) extends CodeBERT by incorporating a data flow graph to capture structural dependencies in code. This enhancement enables the model to better understand variable usage and control flow, making it more effective for bug detection and code summarization tasks. UniXcoder (Guo et al., 2022) unifies code representation across multiple modalities, including text, code, and structured representations. By leveraging cross-modal pretraining, it achieves state-of-the-art results in tasks like code translation and completion.
@@ 2.1 Code Search and Embedding Models
Despite the advancements in code encoding made by Transformer-based models (Xie et al., 2023), they fail to generalize to different tasks and show sub-par performance in basic retrieval tasks due to their lightweight architecture. The semantic gap between natural language and code contributes to the complexity added by the diversity of code representation by programming languages. Utpala et al. (2023) show that previous code embedding models struggle to separate language-specific semantic components from language-agnostic syntactic components during search evaluations. (Utpala et al., 2023) employs methods to eliminate language-specific information, leading to significant performance improvements in retrieval tasks by focusing on language-agnostic components.
@@ 2.2 Benchmarks and Datasets
XLCost (Zhu et al., 2022) is a benchmark dataset created from GeeksForGeeks, utilizing the same function to create parallel code snippets across seven programming languages. CosQA (Huang et al., 2021) is a question-answering dataset with human-annotated query-code pairs, ensuring strong semantic alignment for text-to-code search tasks. While effective for fine-tuning models, its focus on Python limits its applicability to other programming languages.
@@ 2.4 Embeddings and Contrastive Fine-tuning
Improving text embeddings is a key focus, especially for smaller language models. Techniques like contrastive fine-tuning (Khosla et al., 2021) enhance embeddings by aligning semantically similar text pairs, using contrastive and triplet loss (Chopra et al., 2005; Schroff et al., 2015) from Computer Vision. LoRA further refines these embeddings efficiently by introducing low-rank matrices into the attention layers, allowing customized adjustments for specific datasets or tasks Hu et al. (2021) .
@@
The authors of (Galliamov et al., 2024) mention the usage of PEFT methods like LoRA and Prompt Tuning for Text2Code search but do not document any information on the implementation or the analysis of their low-rank adaptation approaches.
Liu et al. (2024) highlight advancements in code retrieval using instruction-tuned large language models (LLMs) and evaluates performance with the NDCG@10 metric across eight tasks. It points out the lack of large, open-source embedding models for code retrieval, as most are proprietary. CodeXEmbed outperforms Voyage Code (AI, 2024) on the CoIR benchmark (Li et al., 2024) by 20% and introduces various retrieval models. However, its large model reliance incurs high storage and computation costs. Our work addresses this by utilizing LoRA adapters on smaller code embedding models, significantly reducing resource overhead while improving performance.
@@ 3 Design and Implementation
Each pretrained model – CodeBERT (Feng et al., 2020), GraphCodeBERT (Guo et al., 2021), and UniXcoder (Guo et al., 2022) – was fine-tuned using the LoRA configuration. For a detailed explanation of the choice of models, see Appendix A1. LoRA employs low-rank decomposition to modify only the attention layers while freezing the remaining weights. This approach reduces memory consumption and accelerates training.
@@
In various experiments conducted with StarCoder (Li et al., 2023), we tested different combinations of the Query, Key, and Value layers within the attention mechanism. This was done to analyze the contributions of the early, middle, and outer layers in low-rank decomposition. For more details about the training procedure, please refer to Appendix A2.
@@ 4.2 Text2Code
To explore the second question highlighted in Section 4, we fine-tune the UniXcoder model for Text2Code retrieval on a single programming language. We use the CosQA (Huang et al., 2021) dataset, as highlighted in Section 4.1. Since CosQA is a question-answering dataset, we experimented with setting the task type for the LoRA Configuration (Houlsby et al., 2019) (Hu et al., 2021) as QUESTION_ANS, which was not feasible for retrieval tasks, since question-answering is a generative task, and would also require the inference using special Roberta (Liu et al., 2019) models encased with question answering, and those suitable for handling parameters like start_position, end_position, etc. The task type was then set to FEATURE_EXTRACTION, the model parameters were frozen, and then the LoRA matrices were trained using Contrastive Trainer. Detailed observations for the same are noted in Appendix A2.
UniXCoder LoRACode - Combined ($r$=64) MRR 31.36 36.02 NDCG@10 35.64 40.44
Figure 2: Increase in Mean Reciprocal Rank and Normalized Discounted Cumulative Gain @ k=10 for LoRACode combined adapter (rank 64) trained on CosQA dataset (Huang et al., 2021), compared to the UniXCoder base model. LoRACode Combined denotes a single adapter fine-tuned over the 3 base models (UniXcoder, GraphCodeBERT, and CodeBERT), but evaluated on UniXcoder.
@@ C PHP Java C++ C# Javascript Python
Table 5: MRR results for Code2Code search of the base models fine-tuned with LoRA modules at ranks 32 and 64, compared with base models. LoRACode Combined denotes a single adapter of given rank fine-tuned over UniXcoder, GraphCodeBERT and CodeBERT), but evaluated on UniXcoder. The average MRR across languages for LoRACode (rank 64) is 46.93 ± 2.28 (95% confidence interval), details of the inference mentioned in Appendix A5.
The LoRA adapters consistently outperformed embedding models like GraphCodeBERT and CodeBERT, demonstrating their efficacy in leveraging low-rank adaptations for code retrieval. Different languages report a significant increase in MRR, ranging from 9.1% for C, 7.47% for C++, 6.54% for Java, 5.82% for C#, 4.43% for Javascript, 3.40% for Python, and 3.8% for PHP. Using a LoRA config with rank 32 also substantially increases MRR over the highest-performing code embedding model UniXcoder (Guo et al., 2022). These findings suggest LoRA’s low-rank decomposition plays a crucial role in improving retrieval accuracy, supporting the use of these adapters for multilingual code search. The LoRA config utilizes only 1.83% to 1.85% of the total trainable parameters, leading to memory-efficient and inexpensive fine-tuning.
A language-wise adaptation for this search task did not equal meaningful results, simply because the XLCoST (Zhu et al., 2022) dataset is designed to perform a cross-lingual code search through parallel samples, using the function_name as the key identifier. See Appendix A4 for detailed observations regarding the time taken to generate embeddings for Code2Code search.
5 Limitations
A key limitation of our approach lies in the narrow domain diversity of available training datasets. CodeSearchNet (Husain et al., 2020) and CosQA (Huang et al., 2021) are primarily built from GitHub code-comment pairs, reflecting open-source conventions and supporting only Text-to-Code retrieval. To enable code-to-code search, we rely on XLCost (Zhu et al., 2022), which provides cross-lingual code pairs. While our experiments demonstrate the effectiveness of LoRA-based fine-tuning, we do not perform a deeper investigation of the exact language-specific representation in the Transformer architecture for these models. Additionally, large-scale LLM-based retrieval was left for future work due to computational constraints.